answer correctly and all parts plz. Thank you :)Freeman Part II: Problem 5 (15 points) Draw the shear and moment diagrams for the beam and determine the maximum shear and th
Answer #1

(5x5) KN 13 KM 80 kn.m K 205m LA205 ht -5m 3 5m RA Steba: calculate Reac Hon force at A and C. EMA =0 -80+ RX10 – 15x5 – 25x7at x= 5m M(5) = 80 + 5.7505) = 108.75 kv.m For section between BC 5<x<10 5 (x+5) KN 80 kn.m 15KN 142 El Efy ol garang Art x15 ton 5kN/m 80kN) 5-75 KN .-34.25kN AV (2) 757 KN с. A x (SFD) -9.25 RN - 34.25 kn/ 1 108.75 (Nik)) AMEA) Not stoought line